OpenStudy (anonymous):

I found out that I have to answer for x so the equation is (x-8)(3x+4)=0

OpenStudy (zehanz):

This equation is of the form a*b = 0. This can be split up as a = 0 or b = 0, So: solve for x: x-8=0 or 3x+4=0

OpenStudy (anonymous):

Okay so then how would I do it?

OpenStudy (zehanz):

What is the solution of x-8=0?

OpenStudy (anonymous):

-8, right

terenzreignz (terenzreignz):

No... You add 8 to both sides of the equation to get x = 8. So 8 is a solution. Now do 3x + 4 = 0

OpenStudy (zehanz):

OK, now the solution of 3x+4 = 0

OpenStudy (anonymous):

8-3x?

OpenStudy (zehanz):

No, if 3x+4=0, you can see that 3x=-4. Why? Because -4+4=0. So the last step is to find the solution of 3x=-4.

OpenStudy (anonymous):

Oh yeah duh, so then it would \[-\frac{ 3 }{ 4}\]

OpenStudy (zehanz):

...I would say it's -4/3

OpenStudy (anonymous):

oh, I always get them backwards. thank you

OpenStudy (zehanz):

YW!